Lewis is offered 20,000.00 as starting salary for a job, with a raise of 2,000.00 at the end of each year of outstanding performance. If he maintain continues outstanding performance, what will his salary be at the end of school year?
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

looks to me like you just add 2,000 at the end of every year he has worked.
the formula of:
for example:
if he starts on january 1, 2020, then
he will make 20,000 in 2020
he will make 22,000 in 2021
he will make 24,000 in 2022
if you want to calculate what he will make in 2021, then 20,000 + (2021 - 2020) * 2,000 = 20,000 + 1 * 2,000 = 22,000
if you want to calculate what he will make in 2022, then 20,000 + (2022 - 2020) = 20,000 + 2 * 2,000 = 24,000
etc.....

Answer by ikleyn(52878)

.

The meaning of your question is DARK and UNCLEAR.